Write a Python script that allows a user to input three exam scores. Display a congratulatory message if the average exam score across all three is greater than 95%. Your solution should follow these general steps;

  1. Obtain the first exam score from the user in the form score/max
    1. Example: 93/100
  2. Obtain the second score
  3. Obtain the third score
  4. Check that the user provided numbers and compute the average test score
  5. Display the average
  6. If the average was greater than 95%, display a congratulatory message
    1. Example: "Great work! Your average was over 95%!!!!"
Answer #1

Screenshot of the code:

Sample Output:

Code to copy:

# Initialize the variable average to 0

average = 0

# Prompt the user to enter the score for 1st exam.

# The score should be in the form score/max

s1=input("Enter exam 1 score: ")

# Prompt the user to enter the score for 2nd exam.

s2=input("Enter exam 2 score: ")

# Prompt the user to enter the score for 3rd exam.

s3=input("Enter exam 3 score: ")

# Extract the marks and convert them to integer format.

# The score is split at the symbol '/' and the part

# which is before '/' is stored in the variable s1.

s1=int(s1.split("/")[0])

# Extract the marks and convert them to integer format.

# The score is split at the symbol '/' and the part

# which is before '/' is stored in the variable s2.

s2=int(s2.split("/")[0])

# Extract the marks and convert them to integer format.

# The score is split at the symbol '/' and the part

# which is before '/' is stored in the variable s3.

s3=int(s3.split("/")[0])

# Compute the average of the scores s1, s2 and s3.

average=(s1+s2+s3)/3

# Print the average computed score.

print("Average = %.1f" %average)

# If the average score is greater than 95, print the

# required statement.

if(average>=95):

   print("Great work! Your average was over 95%!!!!")